If an airplane accelerated down a runway at 3.2 m/s squared for 32.8 seconds until it lifts of the ground what is the distance before takeoff?
1
Expert's answer
2019-09-10T13:48:42-0400

The distance traveled is given by


"X_{f}=X_{o}+V_{ox}*t+\frac{1}{2}a_{x}t^{2}"


Where:

  • Initial position "X_{o}=0m"
  • Final position "X_{f}"
  • initial velocity "V_{ox}=0\frac{m}{s}"
  • Acceleration "a=3.2\frac{m}{s^{2}}"
  • time "t=32.8s"


Numerically evaluating


"X_{f}=0m+0\frac{m}{s}*32.8s+\frac{1}{2}*3.2\frac{m}{s^{2}}*(32.8s)^{2}\\
X_{f}=1721m"


The distance traveled before taking off is

"\boxed{X_{f}=1721m}"